OpenStudy (anonymous):

evaluate: tan(- pi/24)

OpenStudy (freckles):

tan(x) is an odd function so you could look at evaluating -tan(pi/24) instead

OpenStudy (freckles):

\[\frac{\pi}{4}-\frac{\pi}{6}=\frac{6\pi-4\pi}{24}=\frac{2\pi}{24}=\frac{\pi}{12} \\ \frac{\pi}{24}=\frac{1}{2} \cdot \frac{\pi}{12}\]

OpenStudy (freckles):

I don't know if there is a less complicated way but it looks like you can use difference rule then half angle identity